OpenStudy (anonymous):

Calculate the number the possibility of drawing 5 red cards in a row with no replacement

OpenStudy (anonymous):

how many total card are there?

OpenStudy (anonymous):

52 cards in a standard deck. 13 Diamonds and 13 Hearts, all RED.\[\frac{26}{52}\frac{25}{51}\frac{24}{50}\frac{23}{49}\frac{22}{48}=\frac{253}{9996}=0.0253 \]

OpenStudy (anonymous):

that's right